OpenStudy (anonymous):

solving the given equation: -2+4(x-1)=-7-4x -2+4x-4 =-7-4x 6+4x =-7-4x 6-7+4x=7-7+4x I know I made a mistake what do I do next can you explain?

OpenStudy (anonymous):

I believe the answer is (-1/8)=x

OpenStudy (anonymous):

I believe the book gives x=1/8 I need to know the steps

OpenStudy (anonymous):

can you help me

OpenStudy (amistre64):

-2-4+4x = -7-4x -6 + 4x = -7 -4x <-- you have "6+4x". Watch those signs, theyll mess you up everytime. -6 +4x = -7 -4x -6 +7 = -4x -4x 1 = -8x (-1/8) = x

OpenStudy (anonymous):

thank you